OpenStudy (anonymous):

what is the slope of the line that contains point 3, 5 and -1, 7, ? f. -3 g. -1/3 h. -1/4 k 1/3 j 3

OpenStudy (anonymous):

urgent please help

OpenStudy (espex):

\[m=\frac{y_2 - y_1}{x_2-x_1}\]

OpenStudy (anonymous):

???

OpenStudy (espex):

When given two points, you find your slope, m, by taking the difference of y and dividing by difference of x. \[\frac{\delta y }{\delta x} \rightarrow \frac{rise}{run} \rightarrow \frac{y_2 - y_1}{x_2 - x_1}\]